# Ramanujan’s function on small primes

### v2

@@ Abstract.
Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number. For this reason, as an example in a recent article about matrix methods for the study of number-theoretic functions, we studied with computer calculations the function $n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime. We noted that (within the range of our observations) the complex eigenvalues of certain matrices associated to this function exhibit regularities of the following kind: the minimum modulus among the eigenvalues for the matrix associated to $\tau(p_{n})$ appears to oscillate in a nearly periodic fashion with $n$.
@@ 1. Introduction
Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number [Le47] . For this reason, we made computer calculations about the function on positive integers
$\tau_{p}:n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime. Using matrix machinery reproduced in Lemmas 2.1 and 2.2 below, within the range of our observations, the eigenvalues of the matrices $H_{n}$ and $J_{n}$ attached to this function by the two lemmas behave as follows: the minimum modulus among the eigenvalues for each of the two matrices associated to a particular $\tau_{p}(n)$ appear to oscillate with $n$. The connection to Lehmer’s question is that the associated matrix $H_{n}$ has a zero eigenvalue if and only the associated matrix $J_{n}$ does too, and, if they do, $\tau_{p}(n)=0$.
We show plots of the oscillations in Figure 1 and Figure 3 below. We also describe some pleasant patterns among the abscissas of the points on their lower envelopes that seem unlikely to persist unaltered: they seem to suggest that the effect of Ramanujan’s function on the prime sequence is to tune somehow the noise in it, and we do not know how that might work. We have looked at similar compositions of tau with other functions and, for them, this “tuning” is not so clear-cut.
@@ 2. Lemmas
In the context of the theory of symmetric functions, the lemmas in this section are well known. They appear for example, in MacDonald’s book [M]), but the proofs are not easy to locate. (In fact, we have not located them.) We have deposited home-made proofs of these statements in our depository [Bre25]111“Lemmas for ‘Ramanujan’s function on small primes’ ”. In the sequel, we denote the cardinality of a set $S$ by #S\#S and the determinant of a matrix $M$ by $|M|$.
@@
3. Empirical observations about the sequence $\{\tau_{p}(n)\}_{n\geq 1}$
3.1. Characteristic functions of the undeformed matrices of the form $J_{n}(\overline{j_{\tau_{p}}})$ representing the function $\tau_{p}$.
Let $h_{\tau_{p}}(0)=1$ and $h_{\tau_{p}}(n)=\tau_{p}(n)$ for positive integers $n$. Let $j_{\tau_{p}}(n)$ and $h_{\tau_{p}}(n)$ together satisfy equation (D). Given $h_{\tau_{p}}$, the elements of $\overline{j_{\tau_{p}}}=\{j_{\tau_{p}(n)}\}_{n=1,2,...}$ can be computed from equation (D1). By Lemma 2.2 (2), $\tau_{p}(n)=|J_{n}(\overline{j_{\tau_{p}}})|/n!$. Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j_{\tau_{p}}}))(x)=|xI-J_{n}(\overline{j_{\tau_{p}}})|$ be the characteristic polynomial of $J_{n}(\overline{j_{\tau_{p}}})$. The following statements are clearly equivalent: (a) $\tau(p_{n})=0$,
(b) $|J_{n}(\overline{j_{\tau_{p}}})|=0$, and (c) $\Pi_{n}(0)=0$.
(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau_{p}(k).$
In (theoretical) practice, this result would allow us to bypass the construction of the matrices $J_{n}(\overline{j_{\tau_{p}}})$ in our computations, saving time and allowing us to extend our data sets. We have not found occasion to exploit this advantage; if a similar result applied to the $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))(x)$, we would have used it; at present, this is not available to us.
Remark 1.
Let $M$ be an $n\times n$ matrix over a field. For any subset $S\subseteq\{1,2,\ldots,n\}$, let $M[S,S]$ denote the submatrix of $M$ with rows and columns indexed by $S$. The following proposition is well known. (For example, see equation (1.2.13) in Horn and Johnson’s book  [Horn2012].)
@@ 3. Empirical observations about the sequence $\{\tau_{p}(n)\}_{n\geq 1}$
The number of terms in the sum is $\binom{n}{k}$, so the simplest way to reconcile Proposition 1 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau_{p}}})$ somehow force $|J_{n}(\overline{j_{\tau_{p}}})[S,S]|=(-1)^{k}k!\tau_{p}(k)$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$. But this turns out to be too good to be true. [^4] See ‘too good.ipynb’in [Bre25]. Instead, there must be systematic cancellations coming from those symmetries. So far, we have not understood these cancellations.
3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j_{\tau_{p}}})$.
@@
Let $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))=\Pi_{n}^{(1)}(x)$ (say.) (The corresponding objects for $c=0$ exhibit behaviors similar to those we go over here. [^5] See ‘deformed primetau 22oct25.ipynb’ in our repository [Bre25].) Let $V_{n}^{(1)}$ be the set of zeros of $\Pi_{n}^{(1)}(x)$ and let $\mu_{n}^{(1)}=\min_{v\in V_{n}^{(1)}}|v|$; then it appears that the graph of the pairs $(n,\mu_{n}^{(1)})$ lies on an oscillating curve.
Figure 1. Minimum moduli of the roots of the $\Pi^{(1)}_{n}(x)$
and their logarithms.
Figure 2. Minimum moduli of the roots of undeformed $\Pi_{n}(x)$.

### v3

# Ramanujan functions on small primes

@@ Abstract.
Lehmer asked whether Ramanujan’s tau function has any zeros and showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number. Certain matrix identities originating in the theory of symmetric functions express members of arithmetic sequences as determinants. We study empirically the eigenvalues of the matrices underlying them with the goal of estimating how close the eigenvalues of the matrix associated to the number $\tau(p)$ for $p$ prime get to zero, because (with our set-up) a zero eigenvalue would entail that $\tau(p)$ vanishes. We also studied corresponding eigenvalues of several cusp forms coming from elliptic curves.
@@ 1. Introduction
Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number [Le47]. Matrix identities ([M], Lemmas 2.1 and 2.2 below) govern certain additive convolution identities. We study eigenvalues of these matrices empirically, especially eigenvalues associated to the function $n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime, because (with our set-up) a zero eigenvalue would entail that $\tau(p_{n})$ vanishes. After a procedure we will call “deformation”, the minimum modulus among the eigenvalues for a given $n$ appears to oscillate in a nearly periodic fashion with $n$. (We also saw this phenomenon in the sequence $\{\tau(p_{n+1}-p_{n})\}_{n=1,2,...}$; we omit it here.) We have also observed oscillatory behavior in several newforms coming from elliptic curves in virtue of the Modularity Theorem. To encompass those observations, given the appropriate uniformizing function $q_{*}(z)=q_{*}$ (say) from the upper half plane $\mathbb{H}$ to the unit disk for a cusp form $f(z)=\sum_{n}a(n)q_{*}^{n}$, we have referred in our title to the function $n\mapsto a(n)$ as a Ramanujan function.
@@ 2. Lemmas
The lemmas in this section are well known. They appear, for example, in MacDonald’s book [M]), but the proofs are not easy to locate. (In fact, we have not located them.) We have placed home-made proofs of these statements in our repository [Bre25]111“Lemmas for ‘Ramanujan’s function on small primes’ ”.
In the sequel, we denote the cardinality of a set $S$ by #S\#S and the determinant of a matrix $M$ by $|M|$.
@@
Now let $M$ be an $n\times n$ matrix over a field. For any subset $S\subseteq\{1,2,\ldots,n\}$, let $M[S,S]$ denote the submatrix of $M$ with rows and columns indexed by $S$. The following proposition is well known. (For example, see equation (1.2.13) in Horn and Johnson’s book  [Horn2012].)
@@
3. The sequences $\{p_{n}\}_{n\geq 1}$ and $\{\tau(p_{n})\}_{n\geq 1}$
@@
We will freely abuse the notation by employing the $H,h,J$, and $j$ symbols to denote objects in a way that depends on the context without, we hope, confusing anyone.
3.1. The sequence $\{\tau(p)\}_{p\text{ prime}}$ and its developments
Let $p_{n}$ denote the $n^{th}$ prime, $h(0)=1$ and $h(n)=\tau(p_{n})=$ (say) $\tau_{p}(n)$ for positive integers $n$. Let $j(n)$ and $h(n)$ together satisfy equation (D). Given $h$, the elements of $\overline{j}=\{j\}_{n=1,2,...}$ can be computed from equation (D1). By Lemma 2.2 (2), $\tau_{p}(n)=|J_{n}(\overline{j})|/n!$. Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j}))(x)=|xI-J_{n}(\overline{j})|$ be the characteristic polynomial of $J_{n}(\overline{j})$. Plots 3 and 4 show the behavior of the minimum modulus among the eigenvalues of $\Pi^{(1)}_{n}$ and $\Pi_{n}$ respectively. They are relevant to Lehmer’s question because the following statements are equivalent:
(a) $\tau(p_{n})=0$,
(b) $|J_{n}(\overline{j})|=0$, and
(c) $\Pi_{n}(0)=0$.
Remark 1.
(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau_{p}(k).$
This result allows bypass of the construction of the matrices $J_{n}(\overline{j})$, in computations, saving time and allowing the extension our data sets. We have not exploited this advantage; if a similar result applied to the deformations $\chi(J_{n}^{(c)}(\overline{j}))(x)$ (see section 3.2) we would have used it but, at present, this is not available to us.
Remark 2.
The number of terms in the sum in Proposition 2 is $\binom{n}{k}$, so the simplest way to reconcile Proposition 2 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau_{p}}})$ somehow force $|J_{n}(\overline{j_{\tau_{p}}})[S,S]|=(-1)^{k}k!\tau_{p}(k)$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$. But this turns out to be too good to be true. [^4] See ‘too good.ipynb’in [Bre25]. Instead, there must be systematic cancellations coming from those symmetries. So far, we have not understood these cancellations.
3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j})$ representing the sequence $\{\tau(p_{n})\}_{n\geq 1}$
Let $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))=\Pi_{n}^{(1)}(x)$ (say.) (The corresponding objects for $c=0$ exhibit behaviors similar to those we go over here. [^5] See ‘deformed primetau 22oct25.ipynb’ in our repository [Bre25].) Let $V_{n}^{(1)}$ be the set of zeros of $\Pi_{n}^{(1)}(x)$ and let $\mu_{n}^{(1)}=\min_{v\in V_{n}^{(1)}}|v|$; then it appears that the graph of the pairs $(n,\mu_{n}^{(1)})$ lies on an oscillating curve. As a control, we repeat the calculation for $h(n)=\tau(p_{n}+1)$. Figure 2 shows no oscillatory behavior for this choice of $h$, indicating that the choice of prime indices per se was decisive.
@@ 3.1. The sequence $\{\tau(p)\}_{p\text{ prime}}$ and its developments
(For contrast, we reproduce in Figure 4 a plot of the mimimum moduli for each $n$ for the undeformed characteristic polynomials $\Pi_{n}(x)$. We have not attempted the sort of analysis we just discussed for the $\Pi_{n}^{(1)}(x)$.)
Fourier analyses were done by the Anthropic AI “Claude” [Anthropic2026]. (We will supply more details and also put Claude’s analyses verbatim in our depository in later drafts.)
Next, with $h_{n}=\tau_{p}(n)$, we used equation (D1) to compute the sequence $\overline{j_{\tau_{p}}}$. We then “deformed” it by prepending $c$ to $\overline{j_{\tau_{p}}}$. We denote this deformation as $\overline{j^{(c)}_{\tau_{p}}}$. Then, once more applying equation (D1), we compute a sequence $\overline{h^{(c)}_{\tau_{p}}}$ from $\overline{j^{(c)}_{\tau_{p}}}$. We formed the list of pairs $(c,n!h^{(c)}(n))$. (These are just the values of the determinants from Lemma 2.2 in this situation; but it is a situation in which we can bypass calculating the determinants and, by that expedient, achieve significant time savings.)
It seemed likely that these pairs would describe a polynomial in $c$, $D_{n}(c)$ (say.) We found by Lagrange interpolation this is so within the range of our observations, and that the degree of $D_{n}(c)$ appears to be $n$.
We found the minimum modulus among the roots of $D_{n}(c)$ and plotted these minima against $n$. In the case $c\in\mathbb{Z}^{\geq 0}$, the result was the same kind of oscillating wave form, as far as we could see (literally.) [^7] Figure 5, Table 7, and notebook ‘tauprime many c’s 18feb26’ in [Bre25]. The reader will notice that we ran the routines twice, at 64-bit and 100-bit precision; the results were indistinguishable. The Fourier analysis of this waveform is described in Table 2.
4. Oscillatory behavior in newforms from elliptic curves
Because the cusp form $\Delta$ is the generating function of the multiplicative Ramanujan function, which displays oscillatory behavior in the range of our observations, we searched a source of cusp forms that are generating functions of multiplicative functions, and found similar oscillatory behavior: Cremona’s database [LMFDB] of cusp forms $\sum_{n}a(n)q_{*}^{n}$ for an appropriate uniformizing parameter $q_{*}(z)$ associated to elliptic curves in virtue of the Modularity Theorem.
These cusp forms have weight two at various levels. [^8] SageMath notebooks corresponding to Table 4 are in [Bre25]. In the limited range of our observations, some deformed curves (in the sense we have been using that term) show oscillatory waveforms, and some that do not, do show oscillatory behavior when the sequences $\{h\}$ satisfy $h(n)=a(1+p_{n})$. [^9] We had thought, mistakenly, that our code was checking the coefficients $a(p_{n})$. To assure ourselves that our present interpretation of the code is the correct one, we recomputed some of the plots with code that checks it explicitly. For example, see the notebooks “curve 11a1 $h(n)=a(p_{n}+1)$ 8mar26” and “curve 14a1 $h(n)=a(p_{n}+1)$ 9mar26” in [Bre25]. Table 3 summarizes Claude’s findings. [^10] The analysis of the curve $\tau(p_{n}),c=0$ was carried out in the notebook ‘deformed c=0 primetau3feb26’ and that of the curve $\tau(p_{n}),c=1$, was done in the notebook ‘tauprime=h logarithmic envelopes prec100 n400 1dec25’, both stored in [Bre25]. The curve labels indicate characteristics of the associated elliptic curves. (See the explanation in section 6.1 below of Table 8, column A.) The column labeled “cc” indicates the deformation parameter in the sense described above. The column labeled “Indices” distinguishes between matrices for $h(n)=a_{n}$ and $h(n)=a_{p_{n}}$. More detail will follow in subsequent drafts of this article.
5. Questions
We have a dilemma. The plots for the undeformed matrices are a mess, therefore, unpromising as sources of insight about the distances of their eigenvalues from the origin, but their determinants represent the underlying $h$ sequences in a straightforward way. The plots for the deformed matrices seem more promising. It is possible that eventually we will be able to use them to formulate some guesses bounding the plotted minimum modulus eigenvalues away from zero. But, so far, we cannot use such conjectures to bound their plotted minimal modulus eigenvalues in the undeformed setting away from zero, and that is what we need. Beyond the matter of making guesses, we do not know why the oscillations we see in the deformed setting occur, why they are absent in the undeformed setting, or very much about how the two settings are related. Other questions are: Are the members of $\overline{h^{(c)}_{\tau_{p}}}$ Fourier coefficients of a modular form? What conditions on $h(n)$ produce oscillatory behavior among the $h^{(c)}$?
To help with comparisons between the presence or absence of oscillatory behavior and characteristics of the objects that do or do not manifest it, we list cusp forms and associated invariants computed using SageMath native functions [sagemath] (Tables 8 - 11.) The tables are explained in the next section. [^11] See ‘cusp form gp invariants 24feb26’ in [Bre25]. Tables 8 - 11 and the key to them in the next section were created to our specifications by the AI “Claude”[Anthropic2026], and edited by the author.
6. Key to Tables 8 - 11
The tables enumerate several invariants for Ramanujan’s $\Delta$ function and elliptic curves over $\mathbb{Q}$ with conductor $N<54$ and unique Fourier expansions in the Cremona database [6]. (Thus there is only one curve labeled NXm, as explained below, for each NX.) The key omits discussion of the two rightmost columns in each table. These are yes or no responses to the questions stated in the table captions. At the time of the present draft, they are entirely subjective.
6.1. Table 8: Basic Arithmetic Invariants
Column A: Curve Label
The Cremona label uniquely identifies an elliptic curve over $\mathbb{Q}$ in the form $NXm$, where $N$ is the conductor, $X$ is a letter indicating the isogeny class (a, b, c, …), and $m$ is a number ordering curves within the isogeny class by discriminant. Isogenous curves have the same associated cusp form (“newform”).
For Ramanujan’s $\Delta$ function, we use the label “Delta”.
Code:
label = E_cremona.label() # Line 267
# For Delta: label = "Delta" # Line 176
References: The Cremona labeling scheme is documented in [5]. Isogeny class structure is discussed in [13, Chapter VII].
Column B: Minimal Weierstrass Form
The minimal (global) Weierstrass model of the curve, reported as the coefficient vector $a_{1},a_{2},a_{3},a_{4},a_{6}]$ in the equation
$y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{2}x^{2}+a_{4}x+a_{6}.$
Code:
E = EllipticCurve(label) # e.g. EllipticCurve(’37a1’)
ainvs = list(E.ainvs()) # returns [a1, a2, a3, a4, a6]
References: Minimal Weierstrass models: [13, Section VII.1]. Uniqueness: [13, Proposition VIII.8.2].
Column C: Weight
The weight of the associated modular form. Code:
weight = 2 # All elliptic curves, Line 286
# For Delta: weight = 12 # Line 181
References: Modularity Theorem: [20, 17, 1]. Weight of modular forms: [diamond-shurman, Chapter 5].
Column D: Torsion Order
The order $|E(\mathbb{Q})_{\text{tors}}|$ of the finite torsion subgroup of the Mordell-Weil group. Code:
torsion_order = int(E.torsion_order()) # Line 289
# For Delta: torsion_order = "N.A." # Line 178
References: Mazur’s theorem [9]. Computational methods: [3, Algorithm 7.4.12].
Column E: Torsion Structure
The group structure of $E(\mathbb{Q})_{\text{tors}}$ as an abelian group, expressed in the form $\mathbb{Z}/n_{1}\mathbb{Z}\times\cdots\times\mathbb{Z}/n_{k}\mathbb{Z}$ where $n_{i}|n_{i+1}$ (invariant factors).
Code:
def torsion_structure_string(E):
T = E.torsion_subgroup()
inv = T.invariants() # Returns tuple of invariant factors
if len(inv) == 0:
return "trivial"
else:
return "x".join([f"Z/{n}Z" for n in inv])
# Line 101-109
References: The SageMath method torsion_subgroup() implements the algorithm from [18, Section 8.7]. Structure theorem for finitely generated abelian groups: [7, Theorem 5.2]. Computation: [18].
Column F: Rank
The rank $r$ of the Mordell-Weil group $E(\mathbb{Q})$. Code:
try:
rank = int(E.rank())
rank_certain = True
except Exception:
rank = -1
rank_certain = False
# Lines 293-298
References: SageMath’s E.rank() implements the algorithm of [5], which combines 2-descent with verification using the BSD conjecture for rank 0 and 1 cases. Mordell-Weil theorem: [10, 19, 13]. BSD conjecture: [2]. Computational methods: [5].
6.2. Table 9: Mordell-Weil Generators and Galois Representations
Column A: Curve Label
Same as column A of Table 8.
Column B: $j$-Invariant
The $j$-invariant $j(E)\in\mathbb{Q}$.
Code:
E = EllipticCurve(label)
j = E.j_invariant() # returns an element of $Q$
References: $j$-invariant and the $b$-, $c$-invariants: [13, Section III.1]. Relation to CM: [14, Appendix A].
Column C: Generator Canonical Heights
Néron-Tate height $\hat{h}(P)$.
Code:
def generator_heights_string(E):
try:
gens = E.gens()
if len(gens) == 0:
return "none"
heights = [float(g.height()) for g in gens] # CORRECTED
return "|".join([f"{h:.4f}" for h in heights])
except Exception as ex:
return f"error:{ex}"
# Lines 111-119
References: Canonical height theory: [13, Chapter VIII]. Computational methods: [5, 15].
Column D: Regulator
The determinant of the height pairing matrix.
Code:
try:
reg = float(E.regulator())
except Exception:
reg = -1.0
# Lines 306-309
References: Regulator definition: [13, Section X.6]. BSD conjecture: [2, 8].
Column E: Galois Representation Surjective for All Primes?
For a prime $\ell$, this is the Galois representation
$\rho_{E,\ell}:\text{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\text{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$
arising from the action of the absolute Galois group on the $\ell$-torsion $E[\ell]$.
Code:
try:
rho = E.galois_representation()
non_surj = list(rho.non_surjective())
gal_surjective = (len(non_surj) == 0)
except Exception as ex:
gal_surjective = "error"
# Lines 313-317
References: Serre’s open image theorem: [11, 12]. Computational aspects: [16].
6.3. Table 10: Galois Exceptional Primes and Modular Data
Column B: Galois Image Type
A classification of the image of the mod-$\ell$ Galois representation. Possible types: “surjective” ($\text{im}(\rho_{E,\ell})=\text{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ for all $\ell$); “non-surjective at $\ell_{1},\ldots,\ell_{k}]$” (list of exceptional primes); or “CM” (curve has complex multiplication, image contained in normalizer of split/non-split Cartan subgroup). The numbers in brackets are the lists of the finite set of primes $\ell$ for which $\rho_{E,\ell}$ is not surjective. [11, 12, 16] For non-CM curves over $\mathbb{Q}$, this set is always finite (Serre).
Code:
non_surj = list(rho.non_surjective())
gal_exceptional_primes = "|".join([str(p) for p in non_surj]) \
if non_surj else "none"
# Lines 318-319
def galois_image_type(E):
try:
rho = E.galois_representation()
if E.has_cm():
return "CM"
non_surj = rho.non_surjective()
if len(non_surj) == 0:
return "surjective"
else:
return f"non-surjectiveat{list(non_surj)}"
except Exception as ex:
return f"error:{ex}"
# Lines 121-131
References: Image classifications: [11, 12, 16] CM theory: [14, Chapter II].
Column C: Index of $\Gamma_{0}(N)$ in $\text{SL}_{2}(\mathbb{Z})$
(NN is the conductor of the curve.) Code:
def X0N_data(N):
try:
G = Gamma0(N)
idx = G.index()
genus = G.genus()
cusps = G.ncusps()
return idx, genus, cusps
except Exception as ex:
return "error", "error", "error"
# Lines 133-141
References: Congruence subgroups: [diamond-shurman, Chapter 1]. Index formula: [diamond-shurman, Proposition 1.4.4].
6.4. Table 11: Modular Curves and L-functions
Column B: Genus of $X_{0}(N)$
NN is the conductor. $X_{0}(N)$ is the compactification of the quotient $\Gamma_{0}(N)\backslash\mathbb{H}$ where $\mathbb{H}$ is the upper half-plane.
Code:
G = Gamma0(N)
genus = G.genus()
# Line 136
References: Modular curves: [diamond-shurman, Chapter 3]. Genus formula: [diamond-shurman, Proposition 3.8.2].
Column C: Number of Cusps of $X_{0}(N)$
Code:
cusps = G.ncusps()
# Line 137
References: Cusp counting: [diamond-shurman, Proposition 3.8.3].
Column D: Has Complex Multiplication?
Code:
try:
is_cm = E.has_cm()
cm_disc = int(E.cm_discriminant()) if is_cm else 0
except Exception:
is_cm = "error"
cm_disc = "error"
# Lines 326-330
References: CM theory: [14, Chapter II], [4]. The 13 CM $j$-invariants: [14, Appendix A].
Column E: CM Discriminant
If $E$ does not have CM, this field is 0.
Code:
cm_disc = int(E.cm_discriminant()) if is_cm else 0
# Line 328
Column F: Root Number
The root number $w\in\{+1,-1\}$. Code:
try:
root_number = int(E.root_number())
except Exception:
root_number = "error"
# Lines 333-336
References: Functional equation: [18, Chapter 16]. Root number and BSD: [2, 8]. Computational methods: [5, Section 2.14].
Potential issues
  1. (1)
    Rank computation: For curves of large rank ($\geq 2$), E.rank() may be slow or fail. The BSD conjecture is used heuristically for rank 0 and 1. All curves in this dataset ($N\leq 54$) have rank $\leq 1$, so all rank computations are provably correct.
  2. (2)
    Generator computation: For rank 1 curves, E.gens() returns provably correct generators using Cremona’s algorithm. Heights are computed to machine precision ($\approx 10^{-15}$ relative error).
  3. (3)
    Galois representations: The method non_surjective() is provably correct for the primes it returns, but could theoretically miss exceptional primes beyond SageMath’s computational limits. For $N\leq 54$, all computations are complete.
7. Tables